# Solving Football by Exploiting Equilibrium Structure of 2p0s Differential Games with One-Sided Information

### v3

@@ 4 A Primal-Dual Reformulation of the Game
Next we present Thm. 4.2, which proves that $(\{\eta_{i,{\tau}}^{\dagger}\},\zeta_{\tau}^{\dagger})$ computed from $\text{P}_{1}$ and $\text{P}_{2}$ approaches the equilibrium of $G$ when $\tau$ is sufficiently small. The theorem uses P1’s loss in $G$ when they play NE in $G_{\tau}$: $V_{1}(t,x,p):=\max_{\zeta\in\mathcal{Z}}J(t,x,p;\{\eta_{i,{\tau}}^{\dagger}\},\zeta)$, and P2’s loss in $G^{*}$ when they play NE in $G_{\tau}^{*}$: $V_{1}^{*}(t,x,p):=\max_{\{\eta_{i}\}\in\{\mathcal{H}^{i}\}^{I}}J^{*}(t,x,\hat{p};\{\eta_{i}\},\zeta_{\tau}^{\dagger})$.
@@
If A1-5 hold, there exists $C>0$, such that $V_{1}(t,x,p)-V(t,x,p)\in[0,C(T-t)\tau]$ for any $(t,x,p)\in[0,T]\times\mathcal{X}\times\Delta(I)$, and $V_{1}^{*}(t,x,\hat{p})-V^{*}(t,x,\hat{p})\in[0,C(T-t)\tau]$ for any $(t,x,\hat{p})\in[0,T]\times\mathcal{X}\times\mathbb{R}^{I}$.
@@
$V\leq V_{1}$ because P1’s NE in $G_{\tau}$ is not NE in $G$ and $V$ is unique. Compared with $G$, P2 has an advantage in $G_{\tau}$ since he plays best responses to the actions to be played by P1, thus $V_{1}\leq V_{\tau}$. Now we just need to show $\max_{x,p}|V(t,x,p)-V_{\tau}(t,x,p)|\leq C(T-t)\tau$. This is done through a recursion that leverages (1) the consistency property of the Bellman backup (denoted by $T_{\tau}[V]$, $\text{P}_{1}$): $|V(t,x,p)-T_{\tau}[V](t+\tau,x,p)|\leq C\tau^{2}$, i.e., backing up the value via $\text{P}_{1}$ causes a error quadratic in $\tau$, and (2) the boundary condition $V(T,x,p)=V_{\tau}(T,x,p)$. The consistency property is derived from the fact that values of first-order Hamilton-Jacobi equations are Lipschitz and $\omega$-semiconcave under A1-5. The same technique can be applied to $G^{*}_{\tau}$. Full proof in App. C. We note that Cardaliaguet (2009) provided the recursion sketch of the theorem similar to ours without explaining the quadratic error or the consistency property.
@@ 5.1 Value approximation
We discretize time as $\{k\tau\}_{k=0}^{K}$, $\tau=T/K$. Let $\mathcal{S}=\{(x,p)_{i}\}_{i\in[|\mathcal{S}|]}$ be a collocation set. We solve $\text{P}_{1}$ starting from $t=(K-1)\tau$ at all collocation points in $\mathcal{S}$. The resultant nonconvex-nonconcave minimax problems have size $(\mathcal{O}(I(I+d_{u})),\mathcal{O}(Id_{v}))$ and are solved by DS-GDA (Zheng et al., 2023), which guarantees sublinear convergence on nonconvex-nonconcave problems. To generalize value and strategies across $\mathcal{X}\times\Delta(I)$, a value network is trained on the minimax solutions and used to formulate the next round of minimax at $t-\tau$. $G^{*}_{\tau}$ is solved similarly.
@@
A similar result applies to the dual game. Zanette et al. (2019) discussed a linear value approximator that achieves $C=1$. However, their method requires solving a linear program (LP) for every inference $\hat{V}_{t}(x,p)$ if $(x,p)$ does not belong to the training set $\mathcal{S}$. In our context, incorporating their method would require auto-differentiating through the LP solver for each descent and ascent steps in minimax, which turned out to be too expensive. To this end, we introduce a multigrid scheme to reduce the cost for games with a large $K$.
@@ 5.2 Model-free multiagent reinforcement learning
Recent studies (Rudolph et al., 2025; Sokota et al., 2022) showed that policy gradient methods for MARL, such as PPO and MMD, can effectively solve IIEFGs. We show that the atomic structure can be directly applied to this unified model-free framework with minimal code changes, while yielding significant solution improvement for 2p0s1. For CAMS-DRL, the policy network of P1 takes in the infostate $(t,x,p)$ and outputs $I$ logit vectors $\ell_{k}\in\mathbb{R}^{I}$ and $I$ action prototypes $\mu^{k}\in\mathcal{U}$ for $k\in[I]$. Each logit vector $\ell_{k}$ is transformed via softmax to define the behavioral strategy of type-$i$ P1, i.e., the probability $\eta_{i}(\mu^{k};t,x,p)$ of choosing action $\mu^{k}$. The policy network for P2 takes in $(t,x,p)$ and outputs a single logit $\ell\in\mathbb{R}^{I+1}$ and action prototypes $\mu^{k}$ for $k\in[I+1]$. We directly solve the NE of these policy models using PPO and MMD.
@@ 5.3 Model predictive control
When dynamics $f$ is known and an initial state $(t_{0},x_{0},p_{0})$ is given, we can formulate the primal (resp. dual) minimax objective as the sum over $I^{K}$ (resp. $(I+1)^{K}$) game tree paths (Fig. 1a). CAMS-MPC builds the computational graph for the entire tree where the policy networks follow CAMS-DRL, and applies a minimax solver (e.g., DS-GDA) to this differentiable loss. This method is feasible thanks to the atomic structure of NEs and for small $I$. Since P1 in 2p0s1 games reveals their payoff mid-game, the game tree further collapses after information revelation. E.g., Hexner’s game (see below) has a proven game-tree complexity of $I$, where P1 plays a fixed nonrevealing strategy before splitting to $I$ type-dependent revealing strategies. Due to this collapse, modeling strategies using neural networks turns out to be more effective for the convergence to NEs than using infostate-wise parameterization.
@@ 6 Empirical Validation
We introduce Hexner’s game (Hexner, 1979; Ghimire et al., 2024) to compare CAMS variants with baselines on solution quality and computational cost. This game has an analytical NE (see proof in App. F.1). The dynamics is $\dot{x}_{j}=A_{j}x_{j}+B_{j}u_{j}$ for $j=[2]$, where $x_{j}\in\mathcal{X}_{j}$, $u_{j}\in\mathcal{U}_{j}$, and $A_{j}$ and $B_{j}$ are known matrices. The target state of P1 is $z\theta$ where $\theta$ is drawn with distribution $p_{0}$ from $\Theta$, $|\Theta|=I$, and $z\in\mathbb{R}^{d_{x}}$ is fixed and public. The expected payoff to P1 is:
$\small J(\{\eta_{i}\},\zeta)=\mathbb{E}_{i\sim p_{0}}\left[\int_{0}^{T}\|\eta_{i}(t)\|_{R_{1}}^{2}-\|\zeta(t)\|_{R_{2}}^{2}dt\right.+\left.\|x_{1}(T)-z\theta_{i}\|^{2}_{K_{1}}-\|x_{2}(T)-z\theta_{i}\|_{K_{2}}^{2}\Bigg],\right$.
where $R_{1},R_{2},K_{1},K_{2}\succ 0$ are control- and state-penalty matrices. The goal of P1 is to get closer to the target $z\theta$ than P2. To take information advantage, P1 needs to decide when to home-in to and reveal the target. Analytical NE: There exists a critical time $t_{r}:=t_{r}(T,\{A_{j}\},\{B_{j}\},\{R_{j}\},\{K_{j}\})$. If $t_{r}\in(0,T)$, P1 moves towards $\mathbb{E}[\theta]$ as if he does not know the actual target until $t_{r}$ when he fully reveals the target, i.e., value convexification happens at $t_{r}$. If $t_{r}\leq 0$, P1 homes towards the actual target at $t=0$. P2’s NE mirrors P1. See proof in App. F.1.
@@ 6.1 Effect of atomic NE on value approximation
$U$-invariant convergence of CAMS.
We use a normal-form Hexner’s game with $\tau=T$ and a fixed initial state $x_{0}\in\mathcal{X}$ to demonstrate that baseline algs. suffer from increasing discrete action sizes while CAMS does not. The baselines are SOTA normal-form solvers including CFR+ (Tammelin, 2014), MMD (Sokota et al., 2022), JPSPG (Martin & Sandholm, 2024), and a modified CFR-BR (Johanson et al., 2012) (dubbed CFR-BR-Primal), where we focus on converging P1’s strategy and only compute P2’s best responses. Among these, only JPSPG naturally handles continuous action spaces. All baselines except JPSPG are standard implementations in OpenSpiel (Lanctot et al., 2019). The normal-form primal game has a trivial ground-truth strategy where P1 goes directly to his target. For visualization, we use $d_{x}=4$ (position and velocity in 2D). For baselines except JPSPG, we use discrete action sets defined by 4 grid sizes so that $U=|\mathcal{U}_{j}|\in\{16,36,64,144\}$. All algs. terminate when a threshold of $NashConv(\pi_{i})=\max_{\pi_{i}^{\prime}}V_{i}(\pi_{i}^{\prime})-V_{i}(\pi_{i})$ is met. For conciseness, we only consider solving P1’s strategy and thus use P1’s NashConv. We set the threshold to $10^{-3}$ for baselines and use a more stringent threshold of $10^{-5}$ for CAMS. We then use DeepCFR and JPSPG as baselines for a 4-stage game where $T=1$ and $\tau=0.25$. DeepCFRs were run for 1000 CFR iterations (resp. 100) with 10 (resp. 5) traversals for $U=9$ (resp. $16$). The wall-time costs for game solving are 17 hours using CAMS (baseline), 24 hours for JPSPG, 29 hours ($U=9$) and 34 hours ($U=16$) using DeepCFR, all on an A100 GPU. More details on experiment settings can be found in App.  G.3. Furthermore, JPSPG was run for $2\cdot 10^{8}$ iterations, where each iteration consists of solving a game with a random initial state and type, and performing a strategy update. More details in App. G.4.
@@ 6.2 Effect of atomic NE on model-free MARL
For model-free MARL, we compare CAMS-DRL with MMD, PPO and R-NaD with discrete actions of size 100 formed by pairing each of the ten linearly spaced $x$-direction acceleration between (1,1)(-1,1) and $y$-direction acceleration between (4,4)(-4,4). We test the policy convergence in the normal-form game by comparing P1’s learned policy every 8192 steps (corresponds to 1 iteration in Fig. 5)a against the ground truth policy. MMD, PPO, and R-NaD are implemented as in Rudolph et al. (2025). Results in Fig. 5, which uses the same comparison metric as outlined in Sec. 6.1, show that CAMS-DRL approximates NEs accurately, while PPO and MMD fail completely. R-NaD, on the other hand, shows better performance than MMD and PPO with little to no hyperparameter tuning. This contradicts with observations in Rudolph et al. (2025). We conjecture that PG methods such as PPO and MMD have difficulty converging to non-interior solutions, which is the case in 2p0s1 due to the atomic structure of NEs. We performed the same comparison for a 4-stage game, with an action space size of 100 composed by pairing linearly spaced $x$ and $y$ direction accelerations, all between (12,12)(-12,12). Results in Fig. 5)b show significant improvement in solution accuracy when the atomic structure is exploited.
Figure 5: Comparisons b/w CAMS-DRL and standard PG methods on (a) 1-stage and (b) 4-stage games.
@@ 6.3 Effect of atomic NE on MPC
With known dynamics and $I=2$, a 10-stage Hexner’s (primal) game has at most $2^{10}$ rollouts, making exact policy gradient computation feasible. Using DS-GDA, the convergence to the ground-truth NE takes only 5 minutes on a M1 Pro Macbook. See results in App. H.
@@
The independence of the game-tree complexity from the action space allows us to solve games beyond what IIEFG solvers can afford. Here we model an 11-vs-11 American Football play as a short, two-team pursuit–evasion game in a 2D plane. The horizontal axis represents “downfield” progress and the vertical axis is sideline-to-sideline. Each player follows double-integrator kinematics and chooses acceleration at discrete time steps. Physical contact is captured by a smooth “merge” weight that grows as opponents approach, softly blending their velocities and attenuating their ability to apply new acceleration, so tackles emerge continuously rather than through hard impulses. The offense has two types ($I=2$): an “RB power push” that prefers the RB to advance straight upfield, and a “QB throw” that rewards whichever offensive player gets furthest downfield. With loose bounds on velocity and acceleration, the soft tackle dynamics is control-affine, allowing Isaacs’ condition to hold. Together with differentiable rollouts, the game settings satisfy A1-5, leading to atomic structure of the NE. With $K=10$ and known differentiable dynamics, it is then feasible to solve P1’s (and then P2’s) strategy by applying DS-GDA directly to the minimax objective constituting all $I^{K}$ ($(I+1)^{K}$) rollouts. See App. I for detailed game settings. The convergence takes approximately 30 minutes on a M1 Pro Macbook. The resulting plays are summarized in Fig. 1b,c and animated in the github repo. Qualitatively, the results resemble real football tactics: the offense either tries to push through the defense (aka inside zone play), or goes out wide by faking a move (aka waggle play), while the defense, in response, either close-in on the player with the ball possession or stay back to guard. Importantly, our results also show that the offense conceal their intent on their play selection for 0.5 seconds, which is comparable to coaching analyses that estimate roughly a one-second window before the play becomes clearer (Grabowski, 2020).

### v4

@@ 4 A Primal-Dual Reformulation of the Game
Next we present Thm. 4.2, which proves that $(\{\eta_{i,{\tau}}^{\dagger}\},\zeta_{\tau}^{\dagger})$ computed from $\text{P}_{1}$ and $\text{P}_{2}$ approaches the equilibrium of $G$ when $\tau$ is sufficiently small. The theorem uses P1’s loss in $G$ when they adapt $\{\eta_{i,{\tau}}^{\dagger}\}$ to $G$ by playing a constant action within each time interval. This loss is denoted by $V_{1}(t,x,\hat{p}):=\max_{\zeta\in\mathcal{Z}}J(t,x,p;\{\eta_{i,{\tau}}^{\dagger}\},\zeta)$. Similarly, P2’s loss in $G^{*}$ when they adapt $\zeta_{\tau}^{\dagger}$ to $G$ is denoted by $V_{1}^{*}(t,x,p):=\max_{\{\eta_{i}\}\in\{\mathcal{H}^{i}\}^{I}}J^{*}(t,x,\hat{p};\{\eta_{i}\},\zeta_{\tau}^{\dagger})$.
@@
If A1-5 hold, then for any $\epsilon>0$, there exists $\tau^{*}>0$, such that for any $\tau\in(0,\tau^{*}]$ and the corresponding $(\{\eta^{\dagger}_{i,\tau}\},\zeta^{\dagger}_{\tau})$, $V_{1}(t,x,p)-V(t,x,p)\in[0,\epsilon]$ for any $(t,x,p)\in[0,T]\times\mathcal{X}\times\Delta(I)$, and $V_{1}^{*}(t,x,\hat{p})-V^{*}(t,x,\hat{p})\in[0,\epsilon]$ for any $(t,x,\hat{p})\in[0,T]\times\mathcal{X}\times\mathbb{R}^{I}$.
@@
We first note that $V\leq V_{1}$ because P1’s NE in $G_{\tau}$ is not necessarily a NE in $G$, i.e., P1 can get a better value in $G$ using $\{\eta_{i}^{\dagger}\}$ than $\{\eta^{\dagger}_{i,\tau}\}$. We also have, from cardaliaguet2009numerical, the uniform convergence of $V_{\tau}\rightarrow V$ as $\tau\rightarrow 0^{+}$. Then, it remains to show that $V_{1}$ is “not far” from $V_{\tau}$. To do so, we first show that the per-stage value gap between $V_{1}$ and $V_{\tau}$ is within $\mathcal{O}(\tau^{2})$, which accumulates over time into $\mathcal{O}(\tau)$: $V_{1}\leq V_{\tau}+\mathcal{O}(\tau)$. Then as $\tau\rightarrow 0^{+}$, combined with the uniform convergence of $V_{\tau}$, we conclude that the gap $V_{1}-V\in[0,\epsilon]$. This implies that the equilibrium $\{\eta_{i,\tau}^{\dagger}\}$ derived in $G_{\tau}$ is $\epsilon$-optimal in the original game $G$ as long as $\tau$ is small enough. Full proof in App. C. The same techniques can be applied to $G^{*}_{\tau}$.
@@ 5.1 Value approximation
We discretize time as $\{k\tau\}_{k=0}^{K}$, $\tau=T/K$. Let $\mathcal{S}=\{(x,p)_{i}\}_{i\in[|\mathcal{S}|]}$ be a collocation set. We solve $\text{P}_{1}$ starting from $t=(K-1)\tau$ at all collocation points in $\mathcal{S}$. The resultant nonconvex-nonconcave minimax problems have size $(\mathcal{O}(I(I+d_{u})),\mathcal{O}(Id_{v}))$ and are solved by DS-GDA (zheng2023universal), which guarantees sublinear convergence on nonconvex-nonconcave problems. To generalize value and strategies across $\mathcal{X}\times\Delta(I)$, a value network is trained on the minimax solutions and used to formulate the next round of minimax at $t-\tau$. $G^{*}_{\tau}$ is solved similarly.
@@
A similar result applies to the dual game. zanette2019limiting discussed a linear value approximator that achieves $C=1$. However, their method requires solving a linear program (LP) for every inference $\hat{V}_{t}(x,p)$ if $(x,p)$ does not belong to the training set $\mathcal{S}$. In our context, incorporating their method would require auto-differentiating through the LP solver for each descent and ascent steps in minimax, which turned out to be too expensive. To this end, we introduce a multigrid scheme to reduce the cost for games with a large $K$.
@@ 5.2 Model-free multiagent reinforcement learning
Recent studies (rudolph2025reevaluating; sokota2022unified) showed that policy gradient methods for MARL, such as PPO and MMD, can effectively solve IIEFGs with properly tuned hyperparameters. We show that the atomic structure can be directly applied to this unified model-free framework with minimal code changes, while yielding significant solution improvement for 2p0s1. For CAMS-DRL, the policy network of P1 takes in the infostate $(t,x,p)$ and outputs $I$ logit vectors $\ell_{k}\in\mathbb{R}^{I}$ and $I$ action prototypes $u^{k}\in\mathcal{U}$ for $k\in[I]$. Each logit vector $\ell_{k}$ is transformed via softmax to define the behavioral strategy of type-$i$ P1, i.e., the probability $\eta_{i}(u^{k};t,x,p)$ of choosing action $u^{k}$. The policy network for P2 takes in $(t,x,p)$ and outputs a single logit $\ell\in\mathbb{R}^{I+1}$ and action prototypes $\mu^{k}$ for $k\in[I+1]$. We directly solve the NE of these policy models using PPO and MMD.
@@ 5.3 Model predictive control
When dynamics $f$ is known and differentiable, and an initial state $(t_{0},x_{0},p_{0})$ is given, we can formulate the primal (resp. dual) minimax objective as the sum over $I^{K}$ (resp. $(I+1)^{K}$) game tree paths (Fig. 1a). CAMS-MPC builds the computational graph for the entire tree where the policy networks follow CAMS-DRL, and applies DS-GDA to this differentiable loss. This method is feasible thanks to the atomic structure of NEs and for small $I$. Since P1 in 2p0s1 games reveals their payoff mid-game, the game tree further collapses after information revelation. E.g., Hexner’s game (see below) has a proven game-tree complexity of $I$, where P1 plays a fixed nonrevealing strategy before splitting to $I$ type-dependent revealing strategies. Due to this collapse, modeling strategies using neural networks turns out to be more effective for the convergence to NEs than using infostate-wise parameterization.
@@ 6 Empirical Validation
We introduce Hexner’s game (hexner1979differential; ghimire24a) to compare CAMS variants with baselines on solution quality and computational cost. This game has an analytical NE. The dynamics is $\dot{x}_{j}=A_{j}x_{j}+B_{j}u_{j}$ for $j=[2]$, where $x_{j}\in\mathcal{X}_{j}$, $u_{j}\in\mathcal{U}_{j}$, and $A_{j}$ and $B_{j}$ are known matrices. The target state of P1 is $z\theta$ where $\theta$ is drawn with distribution $p_{0}$ from $\Theta$, $|\Theta|=I$, and $z\in\mathbb{R}^{d_{x}}$ is fixed and public. The expected payoff to P1 is:
$J(\{\eta_{i}\},\zeta)=\mathbb{E}_{i\sim p_{0}}\left[\int_{0}^{T}\|\eta_{i}(t)\|_{R_{1}}^{2}-\|\zeta(t)\|_{R_{2}}^{2}dt\right.+\left.\|x_{1}(T)-z\theta_{i}\|^{2}_{K_{1}}-\|x_{2}(T)-z\theta_{i}\|_{K_{2}}^{2}\Bigg],\right$.
where $R_{1},R_{2},K_{1},K_{2}\succ 0$ are control- and state-penalty matrices. The goal of P1 is to get closer to the target $z\theta$ than P2. To take information advantage, P1 needs to decide when to home-in to and reveal the target. Analytical NE: There exists a critical time $t_{r}:=t_{r}(T,\{A_{j}\},\{B_{j}\},\{R_{j}\},\{K_{j}\})$. If $t_{r}\in(0,T)$, P1 moves towards $\mathbb{E}[\theta]$ as if it does not know the actual target until $t_{r}$ when it fully reveals the target, i.e., value convexification happens at $t_{r}$. If $t_{r}\leq 0$, P1 homes towards the actual target at $t=0$. P2’s NE mirrors P1. See proof in App. F.1.
@@
Figure 4: Exploitability comparison in the normal-form Hexner’s game.
@@ 6.1 Effect of atomic NE on value approximation
Action-size-invariant convergence of CAMS.
We use a normal-form Hexner’s game with $\tau=T$ and a fixed initial state $x_{0}\in\mathcal{X}$ to demonstrate that baseline algorithms suffer from increasing discrete action sizes while CAMS does not. The baselines are SOTA normal-form solvers including CFR+ (tammelin2014solving), MMD (sokota2022unified), JPSPG (martin2024joint), and a modified CFR-BR (johanson2012finding) (dubbed CFR-BR-Primal), where we focus on converging P1’s strategy and only compute P2’s best responses. Among these, only JPSPG naturally handles continuous action spaces. All baselines except JPSPG are standard implementations in OpenSpiel (LanctotEtAl2019OpenSpiel). JPSPG implementation details are in App. G.4. The normal-form primal game has a trivial ground-truth strategy where P1 goes directly to its target. For visualization, we use $d_{x}=4$ (position and velocity in 2D). For baselines except JPSPG, we use discrete action sets defined by 4 grid sizes so that $U=|\mathcal{U}_{j}|\in\{16,36,64,144\}$. All algorithms terminate when a threshold of $NashConv(\pi_{i})=\max_{\pi_{i}^{\prime}}V_{i}(\pi_{i}^{\prime})-V_{i}(\pi_{i})$ is met. For conciseness, we only consider solving P1’s strategy and thus use P1’s NashConv. We set the threshold to $10^{-3}$ for baselines and use a more stringent threshold of $10^{-5}$ for CAMS. We then use DeepCFR and JPSPG as baselines for a 4-stage game where $T=1$ and $\tau=0.25$. DeepCFRs were run for 1000 CFR iterations (resp. 100) with 10 (resp. 5) traversals for $U=9$ (resp. $16$). The wall-time costs for game solving are 17 hours using CAMS (baseline), 24 hours for JPSPG, 29 hours ($U=9$) and 34 hours ($U=16$) using DeepCFR, all on an A100 GPU. More details on experiment settings can be found in App.  G.3. Furthermore, JPSPG was run for $2\cdot 10^{8}$ iterations, where each iteration consists of solving a game with a random initial state and type, and performing a strategy update.
@@
With similar or less compute, CAMS achieves significantly better strategies than DeepCFR and JPSPG in the 4-stage game. Sample trajectories for the 4-stage game are visualized in App. G. Fig. 5 compares the ground-truth vs. approximated NEs for 10-stage games with different initial states. While approximation errors exist, CAMS successfully learns the target-concealing behavior of P1. Averaging over 50 trajectories derived from CAMS, P1 conceals the target until $t_{r}=0.60s\pm 0.06s$ (compared to the ground-truth $t_{r}=0.5s$).
@@ 6.2 Effect of atomic NE on model-free MARL
For model-free MARL, we compare CAMS-DRL with MMD, PPO and R-NaD with discrete actions of size 100 formed by pairing each of the ten linearly spaced $x$-direction acceleration between (1,1)(-1,1) and $y$-direction acceleration between (4,4)(-4,4). We test the policy convergence in the normal-form game by comparing P1’s learned policy every 8192 steps (corresponds to 1 iteration in Fig. 6(a)) against the ground truth policy. MMD, PPO, and R-NaD are implemented as in rudolph2025reevaluating. Results in Fig. 6 (a-b), which uses the same comparison metric as outlined in Sec. 6.1, show that CAMS-DRL approximates NEs accurately, while PPO and MMD fail. R-NaD, on the other hand, shows better performance than MMD and PPO with little to no hyperparameter tuning (see App. H). This contradicts with observations in rudolph2025reevaluating. We conjecture that PG methods such as PPO and MMD have difficulty converging to non-interior solutions, which is the case in 2p0s1 due to the atomic structure of NEs. We performed the same comparison for a 4-stage game, with a discrete action size of 100 composed by pairing linearly spaced $x$- and $y$-direction accelerations, all between (12,12)(-12,12). Results in Fig. 6b show significant improvement in solution accuracy when the atomic structure is exploited.
Figure 6: Comparisons b/w CAMS-DRL and standard PG methods on (a) 1-stage and (b) 4-stage games. (c) Exploitability estimate of the strategies in the Football game at different iterations.
@@ 6.3 Effect of atomic NE on MPC
With known differentiable dynamics and $I=2$, a 10-stage Hexner’s (primal) game has at most $2^{10}$ rollouts, making exact policy gradient computation feasible. Using DS-GDA, the convergence to the ground-truth NE takes 5 minutes on a M1 Pro Macbook. See results in App. I.
@@
The independence of the game-tree complexity from the action space allows us to solve games beyond what IIEFG solvers can afford. Here we model an 11-vs-11 American Football play as a short, two-team game in a 2D plane. The horizontal axis represents “downfield” progress and the vertical axis is sideline-to-sideline. Each player follows double-integrator kinematics and chooses acceleration at discrete time steps. Physical contact is captured by a smooth “merge” weight that grows as opponents approach, softly blending their velocities and attenuating their ability to apply new acceleration, so tackles emerge continuously rather than through hard impulses. The offense has two types ($I=2$): an “RB power push” that prefers the RB to advance straight upfield, and a “QB throw” that rewards whichever offensive player gets furthest downfield. With loose bounds on velocity and acceleration, the soft tackle dynamics is control-affine, allowing Isaacs’ condition to hold. Together with differentiable rollouts, the game settings satisfy A1-5, leading to atomic NE. With $K=10$ and known differentiable dynamics, it is feasible to solve P1’s (and then P2’s) strategy by applying DS-GDA directly to the minimax objective constituting all $I^{K}$ ($(I+1)^{K}$) rollouts. See App. J for detailed game settings. The convergence takes 30 minutes on a M1 Pro Macbook. The resulting plays are summarized in Fig. 1b,c and animated in the github repo. Qualitatively, the results resemble real football tactics: the offense either tries to push through the defense (aka inside zone play), or goes out wide by faking a move (aka waggle play), while the defense, in response, either close-in on the player with the ball possession or stay back to guard. Importantly, our results show that the offense conceal their play selection for 0.5 seconds, which is comparable to coaching analyses that estimate roughly a one-second window before the play becomes clear (wake_forest). Quantitatively, we estimate the exploitability of the strategies and plot its convergence in Fig. 6c. We also provide a comparison of resulting trajectories in App. J.1 when each player switches to their respective BR policy to show that the final policies are minimally exploitable.